30 men repair a road in 56 days by working 6 hours daily. In how many days 45 men will repair the same road by working 7 hours daily?

Answer by stanbon(75887)

30 men repair a road in 56 days by working 6 hours daily. In how many days 45 men will repair the same road by working 7 hours daily?
------
d = k/m
Solve for "k"::
56 = k/30
k = 30*56 = 1680
-----
d = 1580/m
----
d = 1580/45
d = 35.11 (days required at 6 hrs per day for 45 men)
----
d' = (6/7)35.11 = 30.095 days required at 7 hrs per day for 45 men)
===========
Cheers,
Stan H.
===========